Project

General

Profile

Actions

2017-07-19-LH2

6th July 2017; 11:00 BST

Phone details

Notes

Actions

  • Action AK follow through de-bugging of FC datalogger.
  • Action CMcW to install NI on MICEnet machine. - Done
  • Action MT/JB to agree and implement testing of quench valves. - In progress
  • Action PW to modify purge sequence limits in agreement with JB. - In Progress
  • Action AK/PW to determine required connections HMI to PLC Ongoing
    • Action CMcW/JB to search archiver for FC radiation shield temp & helium pressure excursions Ongoing
  • Action PF/AK to propose IOC test program for next week. -stands
  • Action AK/JB to arrange further meeting on alarm handler.
  • Action JB/CW/MT to review SH report. Done - single 100mm dia line is sufficient.
  • Action CW to retrieve RP notes on testing on H2 panel. - checked notes were with Kieran - complete. MT to complete document.
  • Action CW update project plan. complete, attached
  • Action CW to contact AN to review wiki and paperwork. Superceded - cover this meeting.

Preliminary H2 tour

Action JB to direct PF to configure paperwork on Wiki. Completion date 2 weeks from tomorrow.

Require: exec summary
test certificates, burst discs, PRV, lift plate, H2 sensor calibration,

global risk assessment - ref to Hazop(s)

Monitoring and archival of data - PF/AK

Current technical

  • FC pressure rating
  • Refurbishment of pipework from turret to gas panel and plenum.
  • Refurbishment of pipework from plenum to roof and vent

Neon Run

Shifters

  • Atex training

Hall access

Timetable

  • Next meeting - Thurs 20th July
  • AoB

Back to LH2:

Back

Files:

Updated by Whyte, Colin almost 6 years ago · 4 revisions